Find the answer to the nearest tenth.
The length of the base of a triangle is 3 times the length of its altitude. Find the length of the base if the area of the triangle is 54 m^2.

Answer by stanbon(75887)

Find the answer to the nearest tenth.
The length of the base of a triangle is 3 times the length of its altitude. Find the length of the base if the area of the triangle is 54 m^2.
-----------------------
Area = (1/2)base*altitude
------
54 = (1/2)(3h)h
-------
108 = 3h^2
h^2 = 36
altitude = 6 m
base = 3h = 18 m
--------------
Cheers,
Stan H.
--------------